Compound Amount


Adrian invests 350,000 at 16% interest compounded quarterly for 7 years. Find the compound amount of Adrian’s investment?


1
Expert's answer
2021-03-16T08:06:03-0400
"A=P(1+\dfrac{r}{n})^{nt}"

Given

"P=350000"

"r=0.16"

"n=4"

"t=7"


"A=350000(1+\dfrac{0.16}{4})^{4(7)}"

"A= 1,049,546.16"